CONSTRAINING THE FLUX OF LOW-ENERGY COSMIC RAYS ACCELERATED BY THE SUPERNOVA REMNANT IC 443

\underline{NICK INDRIOLO}, Department of Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801; GEOFFREY A. BLAKE, Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125; MIWA GOTO, Max Planck Institute for Astronomy, Königstuhl 17, Heidelberg D-69117, Germany; TOMONORI USUDA, Subaru Telescope, Hilo, HI 96720; THOMAS R. GEBALLE, Gemini Observatory, Hilo, HI 96720; TAKESHI OKA, Department of Astronomy & Astrophysics and Department of Chemistry, University of Chicago, Chicago, IL 60637; BENJAMIN J. McCALL, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

It has long been theorized that supernova remnants (SNR) accelerate the majority of Galactic cosmic rays. Observations in the γ-ray, X-ray, and radio regimes support this theory, at least for cosmic rays with energies above a few GeV. However, there is no direct evidence that SNRs accelerate cosmic rays in the MeV–GeV range. These low-energy cosmic rays are of great importance, as they are the primary means by which H_{2} is ionized in the interstellar medium. Collisions between H_{2}^{+} and H_{2} will rapidly form H_{3}^{+}, a molecule which can then be observed to infer the ionization rate of H_{2}. Using the Subaru and Keck telescopes, we have searched for H_{3}^{+} absorption in sight lines which probe molecular material known to be interacting with the SNR IC 443. By computing the ionization rate of H_{2} in these sight lines, we constrain the flux of low-energy cosmic rays generated by this particular supernova remnant.